### v1

# Rotating axionic AdS4 black hole dressed with a scalar field

Abstract
This paper presents a new four-dimensional axionic-charged rotating black hole with a scalar field, which is defined by a structural function coupling the axionic field and a scalar potential. This configuration is characterized by an integration constant and two parameters. The thermodynamic parameters are obtained via the Euclidean procedure, where the validity of the first law of thermodynamics is ensured. With the above information, we can see that this rotating configuration is a good laboratory from the perspective of the holographic superconductor, in which the angular parameter plays an important role.
I Introduction
Scalar fields non-minimally coupled to gravity have been exhaustively studied in the context of black holes (BHs) for nearly fifty years since the seminal work of N. Bocharova, K. Bronnikov, V. Melnikov Bocharova:1970skc and J. Bekenstein Bekenstein:1974sf ; Bekenstein:1975ts , which connected it to the no-hair conjecture Israel:1967wq ; Israel:1967za ; Carter:1971zc ; Hawking:1971tu . Although for the asymptotically flat situation, the model has been subject to debate from a physical point of view, and even the existence of solutions Bronnikov:1978mx ; Xanthopoulos:1992fm ; Klimcik:1993cia , the inclusion of a cosmological constant allows us to explore static BH configurations with a scalar field with other asymptotic behaviors Martinez:2002ru ; Barlow:2005yd ; Martinez:1996gn ; Martinez:2005di ; Dotti:2007cp ; Charmousis:2009cm ; Anabalon:2012tu ; Bardoux:2012tr ; Ayon-Beato:2015ada ; Ayon-Beato:2024xgp ; Ayon-Beato:2023bzp ; Cisterna:2021xxq ; Bravo-Gaete:2021hza . For a scalar field represented by $\phi$, the Lagrangian $\mathcal{L}_{\phi}$ typically takes on the following structure:
$\mathcal{L}_{\phi}=-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{1}{2}\xi R\phi^{2}-U(\phi),$ (1)
where $\xi$ denotes the non-minimal coupling constant, $R$ is the scalar curvature, and $U(\phi)$ denotes the potential depending on $\phi$.
It is interesting to note that BHs with non-minimal coupling and angular momentum have become a nontrivial task due to the complexity of the field equations, which is reflected in the absence of rich literature about it (see Refs. Natsuume:1999at ; Correa:2012rc ; Astorino:2014mda ; Bardoux:2013swa ; Anabalon:2009qt ; Erices:2017izj ; Baake:2020tgk ; Cardenas:2022jtz ). Nevertheless, as shown in Zhao:2013isa ; Zou:2014gla , a spinning configuration can be derived for the three-dimensional scenario where the model is represented through the action
$\mathcal{S}[g_{\mu\nu},\phi]=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}(R-2\Lambda)+\mathcal{L}_{\phi}\right],$ (2)
with an appropriate structure for the potential $U(\phi)$ as well as $\xi=\frac{1}{8}$. Here, the equations of motion with respect to the metric $g_{\mu\nu}$ as well as the scalar field $\phi$ take the structure:
$\mathcal{E}_{\mu\nu}$ ::=:= $G_{\mu\nu}+\Lambda g_{\mu\nu}-T^{\phi}_{\mu\nu}=0,$ (3)
$\mathcal{E}_{\phi}$ ::=:= $\square\phi-\xi R\phi-\frac{dU}{d\phi}=0,$ (4)
where
$T^{\phi}_{\mu\nu}=\nabla_{\mu}\phi\nabla_{\nu}\phi-g_{\mu\nu}\left(\dfrac{1}{2}\nabla_{\sigma}\phi\nabla^{\sigma}\phi+U(\phi)\right)+\xi\left(g_{\mu\nu}\square\phi^{2}-\nabla_{\mu}\nabla_{\nu}\phi^{2}+G_{\mu\nu}\phi^{2}\right).$ (5)
Using this idea raises the natural question of exploring it for the four-dimensional scenario. This is, starting with a metric ansätze of the form
$ds^{2}=-N(r)^{2}f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}(d\theta+N^{\theta}(r)dt)^{2}+r^{2}d\varphi^{2},$ (6)
where $t\in\mathbb{R}$, $r$ is non-negative, while we consider the scalar field as $\phi=\phi(r)$ and set $\xi=\frac{1}{6}$. Further, if we take into account the field equations, which are listed in the Appendix section, we can obtain the expression for the scalar field $\phi$ from the combinations $\mathcal{E}^{t}_{\phantom{a}t}-\mathcal{E}^{r}_{\phantom{a}r}=0=\mathcal{E}^{t}_{\phantom{a}\theta}$. Nevertheless, the combination $\mathcal{E}^{\theta}_{\phantom{a}\theta}-\mathcal{E}^{\varphi}_{\phantom{a}\varphi}=0$ yields
$\frac{1}{12}r^{2}(\phi^{2}-6)\big{(}(N^{\theta})^{\prime}\big{)}^{2}=0,$ (7)
fixing $N^{\theta}(r)=N_{0}$ as an integration constant and, through a coordinate transformation $\tilde{\theta}=\theta+N_{0}t$, the static configuration is naturally recovered.
The exercise aforementioned demonstrates that the prior action proposal (2) is insufficient, thereby redirecting our efforts toward a suitable election for matter sources. One approach is to adopt the concepts from Bardoux:2012tr ; Bardoux:2012aw ; Caldarelli:2013gqa , where two 3-forms that originate from Kalb-Ramond potentials are introduced, allowing us to obtain a planar version of the hairy BH given in Martinez:2002ru ; Martinez:2005di . In addition, it has recently been shown that when the source is composed of two axionic fields that vary linearly with the coordinates of the planar base manifold, charged planar Anti-de Sitter (AdS) BHs can arise Cisterna:2018hzf ; Cisterna:2019uek , enabling the computation of the DC conductivities due to the linear dependence on the coordinates for the axionic fields, providing a momentum dissipation mechanism Andrade:2013gsa ; Donos:2014cya ; Donos:2013eha , and even the potential to explore other contexts (see, for instance, Refs. Ling:2016lis ; Cisterna:2017jmv ; Wang:2018hwg ; Bhatnagar:2017twr ; BravoGaete:2019rci ; Hernandez-Vera:2024zui ). In particular, as we will see in the following lines, a four-dimensional rotating BHs configuration arises when for the action (2) we consider the inclusion of only one axionic field $\psi$ coupled with a function $\varepsilon(\phi)$:
$\mathcal{S}_{\psi}[g_{\mu\nu},\psi]=-\frac{1}{2}\int d^{4}x\sqrt{-g}\,{\varepsilon(\phi)}{\nabla_{\mu}\psi\nabla^{\mu}\psi}.$ (8)
On the other hand, over the past few decades, the study of black holes in AdS spacetime has gained significant interest due to its crucial role in the AdS/Conformal Field Theory (CFT) correspondence Maldacena:1997re ; 2840490 . This duality unveils a fascinating relationship between gravitational theory and a strongly coupled quantum field theory that lives at the conformal boundary of the bulk gravitational spacetime. Thus, it provides a unique approach to studying the non-perturbative regime of quantum field theory systems, where traditional methods often fall short. Examples of these systems are prevalent in the domains of condensed matter physics, in which high-temperature superconductors figure as prototype systems on which duality can be applied. In this regard, the seminal work in holographic superconductivity successfully modeled a high-temperature superconductor using an Abelian-Higgs-like system in a non-back reacted planar Schwarzschild-AdS4 BH background Hartnoll:2008vx . This groundbreaking model elucidated the mechanism of holographic superconductivity by investigating the onset of the superconducting state driven by a $U(1)$ symmetry, breaking, and analyzing the linear response of the system to electric perturbations.
Subsequent studies incorporated an external magnetic field to the superconducting system, enabling a holographic description of the London equation and the Abrikosov vortex lattice, which is characteristic of type-II superconductors Nakano:2008xc ; Hartnoll:2008kx ; Montull:2009fe ; Maeda:2009vf ; Donos:2020viz ; delaCruz-Lopez:2024chw . Notably, the dynamics of these holographic vortex lattices have been explored using both numerical and analytical techniques, uncovering the effect of the external magnetic field on the arrangement and density of vortices Xia:2021jzh ; Herrera-Mendoza:2022whz . Further realizations of holographic superconductors have considered non-commutative BH backgrounds delaCruz-Lopez:2024ync , and the effects of non-relativistic scaling symmetries using BHs with non-standard asymptotic behaviors Bu:2012zzb ; Lu:2013tza ; Natsuume:2018yrg ; Zhao:2013pva . Additionally, holographic superconductivity on isotropic and anisotropic rotating backgrounds has been studied in Herrera-Mendoza:2022whz ; Herrera-Mendoza:2024vfj ; Sonner:2009fk ; Lin:2014tza ; Srivastav:2019ixc , revealing that the BH rotation affects the properties of the superconducting state in such a way that the condensation of the scalar field is suppressed with increasing rotation. In Herrera-Mendoza:2024vfj , a connection was suggested between the rotation of a 5-dimensional anisotropic BH and the damping effects of quasiparticles in a superconducting material.
Taking all of the above into consideration, in the present paper (i) we will show that a four-dimensional rotating configuration arises upon the inclusion of an axionic field, represented by the action (8), in addition to a structured form for the potential $U(\phi)$. We will see that only one integration constant and two parameters determine this configuration. Along with the previously mentioned, via the Euclidean procedure Gibbons:1976ue ; Regge:1974zd the thermodynamic parameters are obtained, where the validity of the first law of thermodynamics is ensured for this spinning hairy BH. With the above information, we will see that (ii) this rotating BH is a good laboratory from the point of view of the holographic superconductor, where the angular parameter plays a significant role.
Two fundamental dimensions underpin the novelty of the current work. Firstly, the inclusion of an axionic field (8) in the toy model (2)
$\mathcal{S}_{\tiny{\mbox{total}}}[g_{\mu\nu},\phi,\psi]=\mathcal{S}+\mathcal{S}_{\psi},$ (9)
[…]
$-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{1}{2}\varepsilon(\phi)\nabla_{\mu}\psi\nabla^{\mu}\psi\equiv-\frac{1}{2}g_{ab}\nabla_{\mu}\Sigma^{a}\nabla^{\mu}\Sigma^{b},\hskip 8.53581pt\Sigma^{a}=(\phi,\psi),$
with the field space metric given by $ds^{2}=d\phi^{2}+\varepsilon(\phi)d\psi^{2}$. The associated scalar curvature $\mathcal{R}$ of this two-dimensional field space geometry reads
$\mathcal{R}=-\dfrac{d}{d\phi}\left(\dfrac{d\varepsilon(\phi)/d\phi}{\varepsilon(\phi)}\right)-\dfrac{1}{2}\left(\dfrac{d\varepsilon(\phi)/d\phi}{\varepsilon(\phi)}\right)^{2},$
which, in general, is not constant. Unlike the typical situation in supergravity models, where scalar manifolds are often realized as symmetric coset spaces with constant curvature Cremmer:1978km ; Andrianopoli:1996cm , the geometry emerging in this case does not correspond to any standard maximally symmetric space, or their noncompact counterparts Freedman:2012zz . This makes the existence of such a configuration particularly intriguing, suggesting a nontrivial interplay between the scalar sector and the underlying dynamics. Secondly, the inclusion of a rotational component offers an intriguing opportunity to explore holographic superconductors within the framework of the AdS/CFT correspondence.
The plan of the work is organized as follows: In Section II, the new axionic rotating BH in four dimensions is presented, while the thermodynamic analysis of this configuration is performed in Section III. In Section IV, from the BH configuration, we examine a holographic superconductor description. In particular, we will focus on the condensation and the electric conductivity. Finally, Section V is devoted to the conclusions and discussions.
II Four-dimensional spinning axionic black holes dressed with a scalar field
Considering the total action (9) in four dimensions with $\xi=\frac{1}{6}$ and the equations of motion
$\mathcal{G}_{\mu\nu}$ ::=:= $\mathcal{E}_{\mu\nu}-T^{\psi}_{\mu\nu}=0,\qquad\mathcal{G}_{\phi}:=\mathcal{E}_{\phi}-\frac{d\varepsilon}{d\phi}{(\nabla_{\mu}\psi\nabla^{\mu}\psi)}=0,\qquad\nabla_{\mu}\left(\varepsilon(\phi)\nabla^{\mu}\psi\right)=0,$ (10)
with the corresponding contribution to the energy-momentum tensor
$T^{\psi}_{\mu\nu}=\varepsilon(\phi)\left(\nabla_{\mu}\psi\nabla_{\nu}\psi-\dfrac{1}{2}g_{\mu\nu}{\nabla_{\sigma}\psi\nabla^{\sigma}\psi}\right),$ (11)
the solution for the metric (6) is given by
$N(r)$ $=$ $1,\qquad f(r)=\left(\dfrac{r}{\ell}\right)^{2}\left(1-\frac{{\ell^{2}}\eta}{\sqrt{\alpha}}N^{\theta}(r)+\ell^{2}(N^{\theta}(r))^{2}\right),$ (12)
${N^{\theta}(r)}$ $=$ ${\frac{\sqrt{\alpha}}{192}\left[\ln\left(\dfrac{2B+r}{r}\right)+\dfrac{2B(r+B)(2B^{2}+2Br-r^{2})}{r^{4}}\right]},$ (13)
together with
[…]
and a negative cosmological constant
$\Lambda=-\frac{3}{\ell^{2}}.$ (15)
Moreover, the configuration is supported by a self-interacting potential $U(\phi)$ and coupling function $\varepsilon(\phi)$ which read:
$U(\phi)$ $={\dfrac{\phi^{4}}{12\ell^{2}}+\dfrac{\sqrt{6}\eta(\phi^{2}+4\phi+6)(\phi^{2}-4\phi+6)\phi}{1152(\phi^{2}-6)}+\dfrac{\eta(\phi^{4}-36)}{2304}\ln\left(\dfrac{\sqrt{6}-\phi}{\sqrt{6}+\phi}\right)}$
$+{U_{\alpha}(\phi)},$ (16)
$\varepsilon(\phi)$ $={\dfrac{\alpha\phi^{4}}{12(\sqrt{6}-\phi)^{5}(\sqrt{6}+\phi)}},$ (17)
where
$U_{\alpha}(\phi)={\dfrac{\alpha(\phi^{4}-36)F(\phi)}{82944}\ln\left(\dfrac{\sqrt{6}-\phi}{\sqrt{6}+\phi}\right)+\dfrac{\alpha\left(\phi^{4}-6\sqrt{6}\phi^{3}+76\phi^{2}-36\sqrt{6}\phi+36\right)(\phi^{2}+6)^{2}\phi^{2}}{18432(\phi-\sqrt{6})^{7}(\phi+\sqrt{6})},}$ (18)
and
${F(\phi)=\dfrac{3}{16}\ln\left(\dfrac{\sqrt{6}-\phi}{\sqrt{6}+\phi}\right)+\dfrac{3\sqrt{6}(\phi^{2}+4\phi+6)(\phi^{2}-4\phi+6)\phi}{4(\phi^{2}-{6})^{2}(\phi^{2}+6)}}.$ (19)
Here, $\eta>0$, $\alpha>0$ are coupling constants, while $B>0$ is the only integration constant present in the solution.
Several key observations can be made regarding the previously presented spinning solution. Firstly, we note that the rotating parameter $\alpha$ takes an important role at the moment to explore this configuration. In fact, at the limit $\alpha\rightarrow 0^{+}$ with $\eta\neq 0$, from equations (13) and (17), the rotating function $N^{\theta}$ as well as the coupling function $\varepsilon$ become zero, while that the potential takes the form (16) with $U_{\alpha}(\phi)=0$. Additionally, we note that the metric function $f$ and the scalar field $\phi$ do not vanish, allowing us to obtain a static configuration where the line element is given by
$ds^{2}=-N(r)^{2}f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}(d\theta^{2}+d\varphi^{2}),$ (20)
together with
$f(r)=\left(\dfrac{r}{\ell}\right)^{2}\left[1-\frac{\ell^{2}\eta}{192}\ln\left(\dfrac{2B+r}{r}\right)-\dfrac{\ell^{2}\eta B(r+B)(2B^{2}+2Br-r^{2})}{96r^{4}}\right],$ (21)
and $N(r)=1$. If, in addition, we now consider $\eta=0$, from Eq. (16) the potential becomes the conformal one ($U(\phi)\propto\phi^{4}$), obtaining the AdS space-time with planar base manifold
$ds^{2}_{\tiny{\mbox{AdS}}}=-\frac{r^{2}}{\ell^{2}}dt^{2}+\frac{\ell^{2}dr^{2}}{r^{2}}+r^{2}(d\theta^{2}+d\varphi^{2}),$ (22)
and allowing us to obtain a stealth configuration
[…]
Here, $T^{\phi}_{\mu\nu}$ and $T^{\psi}_{\mu\nu}$ were given previously in eqs. (5) and (11) respectively. It is interesting to note that eq. (23) have brought a lot of attention during recent years, allowing you to explore new and interesting solutions (see for example Refs. AyonBeato:2004ig ; AyonBeato:2005tu ; Ayon-Beato:2015qfa ; Ayon-Beato:2013bsa ; BravoGaete:2013djh ; Hassaine:2013cma ; Bernardo:2019yxp ; Hassaine:2006gz ; Alvarez:2016qky ; Erices:2024iah ).
With respect to the scalar field $\phi$, as indicated in equation (14), we note that tends to zero for large $r$. Additionally, when we perform an expansion for the coupling function $\varepsilon(\phi)$ and the potential $U(\phi)$, given in eqs. (16)-(19), we obtain
$\varepsilon(\phi)$ $\simeq$ $\frac{\alpha\phi^{4}}{2592}+\frac{\alpha\sqrt{6}\phi^{5}}{3888}+O\left(\phi^{6}\right),$ (24)
$U(\phi)$ $\simeq$ ${\frac{\phi^{4}}{12\ell^{2}}}-\frac{\sqrt{6}\eta{\phi}^{5}}{3240}-{\frac{\alpha\phi^{6}}{46656}}+O\left(\phi^{7}\right),$ (25)
implying that these expressions vanish trivially when $r\rightarrow+\infty$. In relation to the metric function $f$ and $N^{\theta}$ as defined in eqs. (12)-(13), it is worth noting that at infinity
$N^{\theta}(r)$ $\simeq$ ${\frac{\sqrt{\alpha}{B}^{3}}{18{r}^{3}}}+{\frac{\sqrt{\alpha}{B}^{5}}{30{r}^{5}}}+O\left(\frac{1}{r^{6}}\right),$
$f(r)$ $\simeq$ ${\frac{r^{2}}{\ell^{2}}-\frac{B^{3}\eta}{18r}-\frac{B^{5}\eta}{30r^{3}}+O\left(\frac{1}{r^{4}}\right),}$ (26)
showing that this solution is asymptotically AdS in four dimensions.
From the line element (6), the scalar curvature $R$ reads
$R$ $=$ $-\frac{3N^{\prime}f^{\prime}}{N}-\frac{2fN^{\prime\prime}}{N}-f^{\prime\prime}+\frac{r^{2}\big{(}(N^{\theta})^{\prime}\big{)}^{2}}{2N^{2}}-\frac{4fN^{\prime}}{Nr}-\frac{4f^{\prime}}{r}-\frac{2f}{r^{2}}$ (27)
$=$ ${-\frac{12}{\ell^{2}}+\frac{12\eta N^{\theta}}{\sqrt{\alpha}}-12(N^{\theta})^{2}+\frac{8\eta r(N^{\theta})^{\prime}}{\sqrt{\alpha}}-16rN^{\theta}(N^{\theta})^{\prime}+\frac{\eta r^{2}(N^{\theta})^{\prime\prime}}{\sqrt{\alpha}}-\frac{3}{2}r^{2}\big{(}(N^{\theta})^{\prime}\big{)}^{2}}$
-- $2r^{2}N^{\theta}(N^{\theta})^{\prime\prime}.$
Here, for our notations (${}^{\prime})$ denotes the derivative with respect to coordinate $r$. Then, at the moment to evaluate (27) at $f$ and $N^{\theta}$, we get that around $r=0$: [^1] It is worth mentioning that there exists the situation where $r=-2B<0$. However, this case is dismissed due to the constraints imposed by the range of the radial coordinate $r$.
$R\simeq-\frac{\alpha B^{8}}{192r^{8}}-\frac{\alpha B^{7}}{96r^{7}}+O\left(\frac{1}{r^{6}}\right),$ (28)
providing a curvature singularity located at $r_{s}=0$.
Refer to caption Refer to caption
Figure 1: Left Panel: Graphical representation of ${\ell^{2}f(r)}/{r^{2}}$ as a function of the radial coordinate $r$, with $\eta=B=\ell=1$ for our calculations. Here, Case I (red dotted curve), characterized by $\alpha=1$, indicates the presence of a naked singularity ($04-{\ell^{2}\eta^{2}/\alpha}>$ 0). Case II (black curve), corresponds to the extremal case, given by $\alpha=1/4$. Here, ${\ell^{2}f(r^{*})}/{(r^{*})^{2}}=({\ell^{2}f(r^{*})}/{(r^{*})^{2}})^{\prime}=0$ and this condition emerges when $04-{\ell^{2}\eta^{2}/\alpha}=$ 0. Finally, Case III (dashed blue curve and when $\alpha=1/9$) reveals the situation where $04-{\ell^{2}\eta^{2}/\alpha}<$ 0. Right Panel: Graphical representation of the increasing function ${\ell^{2}f(r)}/{r^{2}}$ at the limit $\alpha\rightarrow 0^{+}$, with $\eta=B=\ell=1$, through the black dotted dashed curve. Here, the location of the event horizon is ensured.
On the other hand, to study the existence of event horizon, we consider the function
${\frac{l^{2}f(r)}{r^{2}}=1-\frac{{\ell^{2}}\eta}{\sqrt{\alpha}}N^{\theta}(r)+\ell^{2}(N^{\theta}(r))^{2}},$
where
${\lim_{r\rightarrow 0+}\frac{\ell^{2}f(r)}{r^{2}}=+\infty,\qquad\lim_{r\rightarrow+\infty}\frac{\ell^{2}f(r)}{r^{2}}=1,}$
while that
${\left(\frac{\ell^{2}f(r)}{r^{2}}\right)^{\prime}=(N^{\theta}(r))^{\prime}\ell^{2}\left(2N^{\theta}(r)-\frac{\eta}{\sqrt{\alpha}}\right)=-\left(\frac{\sqrt{\alpha}B^{3}\ell^{2}(r+B)^{2}}{6r^{5}(2B+r)}\right)\left(2N^{\theta}(r)-\frac{\eta}{\sqrt{\alpha}}\right)}.$ (29)
Here, the increasing or decreasing behavior depends on whether the sign of $2N^{\theta}-\eta/\sqrt{\alpha}$ is negative or positive. For the special case $r=r^{*}$ such that ${N^{\theta}(r^{*})=\eta/(2\sqrt{\alpha})}$ , the equation (29) vanishes, and if
${\displaystyle{\left(\frac{\ell^{2}f(r)}{r^{2}}\right)\Big{|}_{r=r^{*}}=1-\frac{\ell^{2}\eta^{2}}{4\alpha}\leq 0,}}$
the existence of horizons is ensured, as is shown in Figure 1 left panel. For the sake of completeness, at the limit $\alpha\rightarrow 0^{+}$, considering now the metric function (21) and
$\frac{\ell^{2}f(r)}{r^{2}}=1-\frac{\ell^{2}\eta}{192}\ln\left(\dfrac{2B+r}{r}\right)-\dfrac{\ell^{2}\eta B(r+B)(2B^{2}+2Br-r^{2})}{96r^{4}},$
we note that
${\lim_{r\rightarrow 0+}\frac{\ell^{2}f(r)}{r^{2}}=-\infty,\qquad\lim_{r\rightarrow+\infty}\frac{\ell^{2}f(r)}{r^{2}}=1},$
while that
${\left(\frac{\ell^{2}f(r)}{r^{2}}\right)^{\prime}=\frac{\eta\ell^{2}B^{3}(r^{2}+2Br+B^{2})}{6r^{5}(2B+r)}>0},$
ensures the existence of the event horizon, as shown in Figure 1 right panel.
Building upon the analysis and framework established for this new four-dimensional axionic spinning BH configuration and the presence of the integration constant $B$ as well as the coupling constants $\alpha$ and $\eta$, in the next section will delve into the thermodynamic quantities.
III Thermodynamics of the rotating black hole solution
The thermodynamic analysis of the previously presented rotating solution (6)-(19) will be conducted using the Euclidean procedure. In this approach, the partition function is equated with the Euclidean path integral, employing the saddle point approximation around the Euclidean continuation of the solution Gibbons:1976ue ; Regge:1974zd . In this Euclidean framework, the Euclidean time $\tau$ is introduced as an imaginary counterpart to the time coordinate $t$, linked through the relation $\tau=it$ from the metric (6). The radial coordinate $r$ spans from the location of, for our case, the outer horizon denoted as $r_{h}$, to +$\infty$+ ∞, while the Euclidean time is made periodic to avoid conical singularities, adhering to the range $0\leq\tau<\beta$, where $\beta=T^{-1}$ and $T$ is the Hawking temperature. With this ansatz, the reduced Euclidean action reads
$I_{E}$ $=$ $\beta\int_{r_{h}}^{+\infty}dr\int_{0}^{\sigma_{\theta}}d\theta\int_{0}^{\sigma_{\varphi}}d\varphi\Big{\{}N\Big{[}\left(1-\frac{\phi^{2}}{6}\right)(rf^{\prime}+f)-\frac{1}{6}r\phi^{\prime}(-rf\phi^{\prime}+r\phi f^{\prime}+4\phi f)$ (30)
-- $\frac{1}{3}\phi r^{2}f\phi^{\prime\prime}+\Lambda r^{2}+U(\phi)r^{2}-\frac{6p^{2}}{r^{4}(\phi^{2}-6)}\Big{]}+N^{\theta}p^{\prime}\Big{\}}$
[…]
where $B_{E}$ is a boundary term constructed such that the Euclidean action has an extremum $\delta I_{E}=0$ and
$p(r)=-\frac{r^{4}\left(\phi(r)^{2}-6\right)N^{\theta}(r)^{\prime}}{12N(r)}.$ (31)
The variation with respects the dynamical fields $N$, $f$, $\phi$, $p$, $N^{\theta}$ and $\psi$ yield the following equations
$E_{N}$ ::=:= $\left(1-\frac{\phi^{2}}{6}\right)(rf^{\prime}+f)-\frac{1}{6}r\phi^{\prime}(-rf\phi^{\prime}+r\phi f^{\prime}+4\phi f)-\frac{1}{3}\phi r^{2}f\phi^{\prime\prime}+\Lambda r^{2}+U(\phi)r^{2}$
$+$ $\frac{1}{2}\varepsilon(\phi)\left(\partial_{\varphi}\psi\right)^{2}-\frac{6p^{2}}{r^{4}(\phi^{2}-6)}=0,$
$E_{f}$ ::=:= ${\frac{1}{6}r\left(2r(\phi^{\prime})^{2}-r\phi\phi^{\prime\prime}\right)N+\frac{1}{6}r\left(r\phi\phi^{\prime}+\phi^{2}-6\right)N^{\prime}=0},$
$E_{\phi}$ ::=:= ${-\frac{1}{3}\phi r^{2}fN^{\prime\prime}-\frac{1}{6}r\left(6rf\phi^{\prime}+4\phi f+3\phi rf^{\prime}\right)N^{\prime}-\frac{1}{6}\Big{(}4\phi rf^{\prime}+2\phi f+r^{2}\phi f^{\prime\prime}+6\phi^{\prime}r^{2}f^{\prime}}$
$+$ ${12rf\phi^{\prime}+6r^{2}f\phi^{\prime\prime}-6r^{2}U_{\phi}-3\varepsilon_{\phi}\left(\partial_{\varphi}\psi\right)^{2}+\frac{12\phi p^{2}}{r^{4}(\phi^{2}-6)^{2}}\Big{)}N=0},$
$E_{p}$ ::=:= $12Np+r^{4}(\phi^{2}-6)(N^{\theta})^{\prime}=0,$
$E_{N^{\theta}}$ ::=:= $p^{\prime}=0,$
$E_{\varphi}$ ::=:= $\partial_{\varphi}^{2}\psi=0.$
[…]
$\delta B_{E}$ $=$ $\beta\sigma\Big{[}N\left(\frac{1}{6}r^{2}\phi\phi^{\prime}+r\left(\frac{1}{6}\phi^{2}-1\right)\right)\delta f-\frac{1}{6}r^{2}\left(N\left(4f\phi^{\prime}+\phi f^{\prime}\right)+2N^{\prime}\phi f\right)\delta\phi$
$+$ $\frac{N}{3}\phi r^{2}f\delta(\phi^{\prime})-N^{\theta}dp\Big{]}_{r_{h}}^{+\infty}-\beta\int_{r_{h}}^{+\infty}N\varepsilon dr\left[\int_{0}^{\sigma_{\theta}}d\theta\partial_{\varphi}\psi\delta\psi\right]_{0}^{\sigma_{\varphi}},$
where the coordinates $\theta$ and $\varphi$ are assumed in the intervals $0\leq\theta\leq\sigma_{\theta}$ and $0\leq\varphi\leq\sigma_{\varphi}$, while $\sigma$ represents
[…]
Now we are in a position to compute the boundary term $B_{E}$. At the infinity, from eqs. (12)-(13) we have that
$\delta N^{\theta}\big{|}_{+\infty}$ $=$ ${\frac{\sqrt{\alpha}{B}^{2}dB}{6{r}^{3}}}+{\frac{\sqrt{\alpha}{B}^{4}dB}{6{r}^{5}}}+O\left(\frac{1}{r^{6}}\right),$
$\delta f\big{|}_{+\infty}$ $=$ ${-\frac{\eta B^{2}dB}{6r}-\frac{\eta B^{4}dB}{6r^{3}}+O\left(\frac{1}{r^{4}}\right),}$ (32)
while that with respect to $p$, and the scalar fields ($\phi,\psi$):
$\delta\phi\big{|}_{+\infty}$ $=$ $\frac{\sqrt{6}dB}{r}-\frac{2\sqrt{6}BdB}{r^{2}}+O\left(\frac{1}{r^{3}}\right),$
$\delta\psi\big{|}_{+\infty}$ $=$ $\varphi\delta B,\qquad\delta p\big{|}_{+\infty}=\delta\left(-\frac{1}{12}{\sqrt{\alpha}B^{3}}\right).$ (33)
With all these ingredients, we obtain that at the infinite
${\delta B_{E}\big{|}_{+\infty}=\frac{\beta\sigma\eta B^{2}dB}{6}\Rightarrow B_{E}\big{|}_{+\infty}=\frac{\beta\sigma\eta B^{3}}{18}.}$
On the other hand, considering at the horizon $r_{h}$
$\delta\phi\big{|}_{r_{h}}$ $=$ $\delta\phi(r_{h})-\phi^{\prime}\big{|}_{r_{h}}\delta r_{h},\qquad\delta\psi\big{|}_{r_{h}}=\varphi\delta B,$
$\delta f\big{|}_{r_{h}}$ $=$ $-\frac{4\pi}{\beta}\delta r_{h},\qquad\delta p\big{|}_{r_{h}}=\delta\left(-\frac{1}{12}{\sqrt{\alpha}B^{3}}\right),$ (34)
we have
$\delta B_{E}\big{|}_{r_{h}}=\delta\left(\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(rh+B)^{2}}\right)+\beta\Omega\delta(\sigma p)+\beta\Psi\delta(\sigma B)\Rightarrow$
$B_{E}\big{|}_{r_{h}}=\left(\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(rh+B)^{2}}\right)+\beta\Omega(\sigma p)+\beta\Psi(\sigma B),$
where
$p=-\frac{1}{12}{\sqrt{\alpha}B^{3}},$
while $\Omega$ and $\Psi_{a}$ are the chemical potentials corresponding to the angular momentum ${\cal{J}}$ and the axionic charge ${\omega}_{a}$, identified as
$\Omega$ $=$ $\lim_{r\rightarrow\infty}N^{\theta}(r)-N^{\theta}(r_{h})=-N^{\theta}(r_{h}),$ (35)
[…]
$=$ $\frac{\sqrt{\alpha}B^{2}}{12}N^{\theta}(r_{h}).$
With all the above, the boundary term $B_{E}$ is found to be
$B_{E}$ $=$ ${B_{E}\big{|}_{+\infty}-B_{E}\big{|}_{r_{h}}=\frac{\beta\sigma\eta B^{3}}{18}-\left(\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(rh+B)^{2}}\right)-\beta\Omega(\sigma p)-\beta\Psi_{a}(\sigma B).}$ (37)
The identification of the thermodynamic parameters is achieved recalling that the Euclidean action is related to the Gibbs free energy $G$ as
${I_{E}=\beta G=\beta{\cal{M}}-{\cal{S}}-\beta\Omega{\cal{J}}-\beta\Psi_{a}\omega_{a}},$ (38)
where ${\cal{M}}$ is the mass, ${\cal{S}}$ the entropy and, as before, $\Omega$ (respectively $\Psi_{a}$) is the chemical potential associated with the angular momentum ${\cal{J}}$ (respectively axionic charge $\mathcal{\omega}_{a}$). Hence, comparing (37) with (38) one obtains that the entropy reads
${\cal{S}}=\frac{2\sigma\pi r_{h}^{3}(2B+r_{h})}{(rh+B)^{2}},$ (39)
while the mass ${\cal{M}}$, angular momentum ${\cal{J}}$, angular velocity $\Omega$, axionic potential $\Psi_{a}$, and axionic charge ${\cal{\omega}}_{a}$ are given by
${\cal{M}}$ $=$ ${\frac{\sigma\eta B^{3}}{18}},$ (40)
${\cal{J}}$ $=$ ${-\frac{\sigma}{12}{\sqrt{\alpha}B^{3}}},\qquad\Omega=-N^{\theta}(r_{h}),$ (41)
${\Psi_{a}}$ $=$ $\frac{B^{2}\alpha(r_{h}+B)(-r_{h}^{2}+2Br_{h}+2B^{2})}{1152r_{h}^{4}}+\frac{B\alpha}{2304}\ln\left(\frac{2B+r_{h}}{r_{h}}\right)=\frac{\sqrt{\alpha}B^{2}}{12}N^{\theta}(r_{h}),$ (42)
${\omega_{a}}$ $=$ $\sigma B.$ (43)
It is reassuring to check that with the Hawking temperature,
$T=\frac{N(r)f^{\prime}(r)}{4\pi}\Big{|}_{r=r_{h}}={\frac{(N^{\theta}(r))^{\prime}(2\sqrt{\alpha}N^{\theta}(r)-\eta)r^{2}}{4\pi\sqrt{\alpha}}\Big{|}_{r=r_{h}},}$ (44)
the first law
${\delta{\cal{M}}=T\delta{\cal{S}}+\Omega\delta{\cal{J}}+\Psi_{a}\delta{\omega_{a}}},$ (45)
is satisfied.
For the non-spinning situation, represented via the line element (20) with $N=1$, together with the metric function $f$ (21), the scalar field $\phi$ (14) and the potential $U(\phi)$ (16) with $U_{\alpha}(\phi)=0$, we note that the entropy ${\cal{S}}$ as well as the mass ${\cal{M}}$ take the form given in (39) and (40) respectively, while that the Hawking Temperature $T$ reads
$T=\frac{\eta B^{3}(r_{h}+B)^{2}}{24\pi r_{h}^{3}(2B+r_{h})},$
and the first law $\delta{\cal{M}}=T\delta{\cal{S}}$ holds.
IV Holographic superconductor
In this section, we are interested in exploring a holographic superconductor description of the preceding BH configuration. Given this, we find it convenient to focus on a particular case of this background in which the transformation $u\to r_{h}/r$ makes it feasible to locate the outer horizon at $u=1$ and the asymptotic boundary at $u=0$. This is possible, provided we define
${\eta=\left(\dfrac{384}{\ln 3}\right)\dfrac{1}{\ell^{2}}+\left(\dfrac{\ln 3}{384}\right)\alpha},\qquad B=(\sqrt{3}-1)r_{h},$ (46)
which preserves the genuine nature of the parameter $\alpha$ as a rotation parameter. In this manner, by evaluating the above expressions into the background functions (12)-(19) and taking the limit $\alpha\to 0^{+}$, the angular momentum of the background is turned off and a static AdS4 BH configuration is obtained (see eqs. (20)-(21)). Moreover, in the setting (46), the Hawking temperature (44) adopts the simpler form
$T=\dfrac{(3-\sqrt{3})(147456-\ell^{2}(\sqrt{\alpha}\ln 3)^{2})r_{h}}{221184\,\pi\,(\ln 3)\,\ell^{2}}.$ (47)
With this in mind, to construct the superconductor, we consider an additional matter contribution consisting of a Maxwell-complex scalar field system of the form Gubser:2008px ; Hartnoll:2008vx ; Hartnoll:2008kx
$S_{\text{M}}=-\frac{1}{q^{2}}\int{d^{4}x\sqrt{-g}\Biggl{(}\dfrac{1}{4}F_{\mu\nu}F^{\mu\nu}+\dfrac{1}{\ell^{2}}\left(|D_{\mu}\Psi|^{2}+m^{2}|\Psi|^{2}\right)\Biggr{)}},$ (48)
in which $F_{\mu\nu}=2\partial_{[\mu}A_{\nu]}$, while $\Psi$ denotes a complex scalar field characterized by mass $m$ and charge $q$, minimally coupled to the gauge field $A_{\mu}$ through the extended covariant derivative $D_{\mu}=\nabla_{\mu}-iA_{\mu}$. Here, we work in the probe limit $q\gg 1$, where the backreaction of the matter contribution (48) onto the background fields can be effectively ignored. Thus, the decoupled system of field equations is given by
$\nabla_{\mu}F^{\mu\nu}-\dfrac{1}{\ell^{2}}\left(2A^{\nu}|\Psi|^{2}-i\Psi\nabla^{\nu}\overline{\Psi}+i\overline{\Psi}\nabla^{\nu}\Psi\right)$ $=0,$ (49a)
$\nabla^{\mu}\nabla_{\mu}\Psi-2iA^{\mu}\nabla_{\mu}\Psi-i\Psi\nabla_{\mu}A^{\mu}-\left(m^{2}+A_{\mu}A^{\mu}\right)\Psi$ $=0,$ (49b)
which is intended to be solved on the fixed background (6) taking into account eq. (46).
IV.1 Condensation of the scalar field and the electric conductivity
In the Abelian Higgs-like description of holographic superconductors as outlined in eq. (48), the mechanism responsible for the onset of superconductivity is the spontaneous breaking of the $U(1)$ Abelian gauge symmetry. This symmetry-breaking paves the way for a non-zero profile of the scalar field. Consequently, in this context, the scalar field is identified as the order parameter that undergoes condensation below some critical temperature. Here, we would like to investigate whether the scalar field $\Psi$ possesses such a property. To do so, using the gauge symmetry of (48), it is sufficient to consider a real scalar field and a Maxwell field with the simplest dependence
$\Psi=\Psi(u),\qquad{A=A_{t}(u)dt+A_{\theta}(u)d\theta}.$ (50)
In this setting, the system of equations (49a)-(49b) reduces to
$A_{t}^{\prime\prime}+\ell^{2}\dfrac{N^{\theta}({N^{\theta}})^{\prime}}{f}A_{t}^{\prime}+\left[\dfrac{N^{\theta}f^{\prime}}{f}-\dfrac{({N^{\theta}})^{\prime}}{f}\left(\ell^{2}({N^{\theta}})^{2}+f\right)\right]A_{\theta}^{\prime}-\left(\dfrac{2\psi^{2}}{u^{2}f}\right)A_{t}=0,$ (51a)
$A_{\theta}^{\prime\prime}+\left(\dfrac{f^{\prime}}{f}-\ell^{2}\dfrac{N^{\theta}({N^{\theta}})^{\prime}}{f}\right)A_{\theta}^{\prime}+\ell^{2}\dfrac{({N^{\theta}})^{\prime}}{f}A_{t}^{\prime}-\left(\dfrac{2\psi^{2}}{u^{2}f}\right)A_{\theta}=0,$ (51b)
$\Psi^{\prime\prime}+\left(\dfrac{f^{\prime}}{f}-\dfrac{2}{u}\right)\Psi^{\prime}-\dfrac{\ell^{2}}{u^{2}r_{h}^{2}f}\Biggl{(}m^{2}r_{h}^{2}+u^{2}A_{\theta}^{2}-\ell^{2}u^{2}\dfrac{\left(A_{t}-N^{\theta}A_{\theta}\right)^{2}}{f}\Biggr{)}\Psi=0,$ (51c)
where now the primes $(^{\prime})$ stand for derivatives with respect to $u$. In solving this system of equations, we need to specify appropriate boundary conditions at the horizon ($u=1$) and at the conformal boundary of the background ($u=0$). Concretely, at the horizon, we require that the scalar field is finite $\Psi(1)<\infty$, while the components $A_{t}(1)$ and $A_{\theta}(1)$ vanish for $A_{\mu}$ to have a finite norm everywhere. On the other hand, at the boundary, the fields behave as
$\Psi=\Psi_{1}u^{\Delta_{1}}+\Psi_{2}u^{\Delta_{2}},\quad A_{t}=\mu-\rho\left(\dfrac{u}{r_{h}}\right),\quad A_{\theta}=\mu_{\theta}-\mathcal{J}_{\theta}\left(\dfrac{u}{r_{h}}\right),$ (52)
with
$\Delta_{1,2}=\dfrac{1}{2}\left(3\pm\sqrt{9+4m^{2}\ell^{2}}\right).$ (53)
From the holographic standpoint, the coefficient $\Psi_{1}$ (respectively $\Psi_{2}$) is identified as the order parameter $\mathcal{O}$ with scaling dimension $\Delta_{1}$ (respectively $\Delta_{2}$). However, to have well-defined normalizable modes at the Ultraviolet (UV) regime (boundary), we must restrict the scalar field mass to be bounded below as $m^{2}\ell^{2}\geq-\frac{9}{4}$, which defines the Breitenlohner-Freedman (BF) bound for a massive scalar field in AdS4 Breitenlohner:1982bm ; Breitenlohner:1982jf . Besides, $\mu$ and $\rho$ are respectively interpreted as the chemical potential and the charge density operator. A similar interpretation is given to the coefficients of $A_{\theta}$, where $\mu_{\theta}$ stands for a one-dimensional potential vector, source of the density of current operator $\mathcal{J_{\theta}}$.
Using the holographic dictionary, by numerically solving the system (51), we can determine the condensation profile of the order parameter as a function of the temperature. For our analysis, we adopt the canonical Ensemble, maintaining a constant charge density $\rho$. In addition, we fix the mass to $m^{2}\ell^{2}=-2$, a value consistent with the BF bound and that gives the simplest asymptotic structure for the scalar field. This choice sets the scaling dimensions $\Delta_{1}=1$ and $\Delta_{2}=2$ for the dual operators $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$, respectively. Fig. 2 depicts the effect of the BH rotation on the condensation curve for $\mathcal{O}_{1}$ (left panel) and $\mathcal{O}_{2}$ (right panel). As shown, increasing the BH rotation (angular momentum) leads to a reduction in the amplitude of the condensation operator. This behavior aligns with findings from previous holographic models that also incorporates rotating BH backgrounds albeit with a different asymptotic nature Lin:2014tza ; Herrera-Mendoza:2024vfj ; Srivastav:2019ixc .
Figure 2: The condensation profiles for the operators $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ as a function of temperature, considering different values of the rotation parameter $\alpha$. The continuous (black) curve corresponds to the static case, while the dashed (violet), dot-dashed (blue), and dotted (red) correspond to situations with non-zero rotation.
In addition to the above, we present the electric conductivity $\sigma_{\tiny{\mbox{DC}}}$, where the idea is to analyze the linear response of the system to a small perturbation of external sources. Concretely, via fluctuations of the vector potentials $A_{\theta}$ and $A_{\varphi}$, denoted as $\delta A_{\theta}$ and $\delta A_{\varphi}$ on the boundary, and the retarded Green’s function. For this, we need to employ the fluctuations (see Refs. Melnikov:2012tb ; Hartnoll:2008vx ):
$\delta A_{\theta,\varphi}=A_{\theta,\varphi}(r)e^{-i\omega\,t},\quad\delta\phi=\delta\psi=0,$ (54)
where $A_{\theta,\varphi}$ has been introduced to jointly represent the components $A_{\theta}$ and $A_{\varphi}$, while that $\omega$ corresponds to the frequency. Considering the complex fluctuation $A=A_{\theta}d\theta+iA_{\varphi}d\varphi$ on the field equations $\nabla_{\mu}F^{\mu\nu}=0$, the equations of motion read
$A{{}^{\prime\prime}}_{\theta,\varphi}+\frac{1}{2}\left(\frac{f{{}^{\prime}}}{f}+\frac{h{{}^{\prime}}}{h}\right)A{{}^{\prime}}_{\theta,\varphi}+\frac{\omega^{2}}{fh}A_{\theta,\varphi}=0,$ (55)
with
$h(r)=f(r)N(r)^{2}-r^{2}\big{(}N^{\theta}(r)\big{)}^{2}.$
Then, to extract $\sigma_{\tiny{\mbox{DC}}}$, which can be calculated through linear response techniques Hartnoll:2008vx ; Hartnoll:2008kx , given an external oscillating electric field $E_{i}(\omega)$ generating a corresponding current $J_{i}(\omega)$, with $i=\{\theta,\varphi\}$, the electrical conductivity is defined as the ratio
$\sigma_{\tiny{\mbox{DC}}}(\omega)=\frac{J_{\theta,\varphi}(\omega)}{E_{\theta,\varphi}(\omega)},$ (56)
where we can compute the electric conductivity through the Kubo formula as a two-point function Baggioli:2016rdj :
$\sigma_{\tiny{\mbox{DC}}}(\omega)=-\frac{i}{\omega}G^{R}_{J_{\theta,\varphi}J_{\theta,\varphi}}=\frac{A{{}^{\prime}}_{\theta,\varphi}}{i\omega A_{\theta,\varphi}}.$ (57)
Figure 3: The $\sigma_{\tiny{\mbox{DC}}}$ curve for the real (left panel) and imaginary (right panel) parts of the electrical conductivity for different values of the rotation parameter $\alpha$. The continuous (black) curve corresponds to the static case, while the other values correspond to situations with non-zero angular momentum.
Figure 3 shows an example of quantum holographic of electronic wave functions, which appears when rotation parameter $\alpha$ increases. This behavior shows us a macroscopic superconducting condensate phase Santos:2024cwf ; Hovhannisyan:2022gkp . We note that the situation $\omega=0$ represents a divergence characterizing a pole in the imaginary part of the conductivity Baggioli:2016rdj . The breaking of symmetry leads to the non-conservation of momentum. This way, the charges dissipate their momentum, and the current will not persist forever. Thus, a physical consequence of these results for real materials is the appearance of impurities, lattice, and disorder, due to the breakdown of momentum conservation Amoretti:2019cef ; Ammon:2019wci ; Andrade:2017lsc ; Tanhayi:2016uui .
V Conclusions and discussions
In the present work, we construct an axionically charged spinning black hole in four dimensions with a non-minimally coupled scalar field. The above is possible via the inclusion of a suitable potential as well as a structural function coupling to the axionic field. It is interesting to note that, although the metric function $f$ and the rotating function $N^{\theta}$ enjoy logarithmic behavior, this configuration is asymptotically AdS.
As can be seen in Fig. 1 (left panel), this configuration is characterized by two parameters ($\alpha$ and $\eta$) and only one integration constant $B$, allowing us to obtain one or even two horizons. Furthermore, we find that the metric function $f$ and the scalar field $\phi$ do not vanish for $\alpha\rightarrow 0^{+}$ and $\eta\neq 0$. This allows us to obtain a new static hairy BH configuration with one horizon (see Fig. 1 (right panel)).
Additionally, due to the inclusion of the scalar and axionic field, this model allows us the emergence of new thermodynamic parameters through the Euclidean action (30), these are the mass $\mathcal{M}$, Hawking temperature $T$, angular momentum $\mathcal{J}$, angular velocity $\Omega$, as well as the axionic potential $\psi_{a}$, and axionic charge $\omega_{a}$, satisfying first law (45) and the four-dimensional Smarr relation Smarr
$\mathcal{M}=\frac{2}{3}T\mathcal{S}+\Omega\mathcal{J}+\frac{1}{3}\psi_{a}\omega_{a}.$ (58)
In exploring the holographic superconductor description of these spinning BHs, we employed a numerical approach to investigate the condensation of the order parameter as a function of temperature and the rotation parameter (see Fig. 2). Notably, the logarithmic falloff of the blackening function does not alter the typical profile of the condensate. Nevertheless, our results reveal that the BH rotation impacts the order parameter. Specifically, we found that increasing the rotation diminishes the amplitude of the condensation curve, suggesting that, for the chosen parameter values, the presence of rotation suppresses the superconducting state in our holographic system. This effect becomes relevant only at sufficiently high values of the rotation parameter. Together with the aforementioned, the electric conductivity as a function of the frequency $\omega$ is calculated numerically from the spinning configuration through the response of the system to small external sources. This is followed by the calculation of the Green’s functions for the electric currents and the application of the Kubo formula. In this framework, the presence of the axionic field in breaking translational symmetry is crucial for modeling realistic condensed matter systems with momentum dissipation. In our case, the $\sigma_{\tiny{\mbox{DC}}}$ is influenced at the moment to consider the effects of rotation, represented through the parameter $\alpha$, as shown in Fig. 3.
Some interesting directions for future research include, for example, the extension of spinning configurations following this toy model in higher dimensions and exploring a broader range of values for the non-minimal coupling constant $\xi$. Additionally, we can obtain a modification of the first law of BH thermodynamics, considering now the cosmological constant $\Lambda$ as a thermodynamic parameter related to the pressure $P$ of the system Kastor:2009wy ; Kubiznak:2014zwa ; Kubiznak:2016qmn ; Dolan:2010ha ; Cvetic:2010jb , where now the first law takes the form
${\delta{\cal{M}}=T\delta{\cal{S}}+V\delta P+\Omega\delta{\cal{J}}+\Psi_{a}\delta{\omega_{a}}}.$
Here, $V$ corresponds to the conjugate thermodynamic volume and $\cal{M}$ now takes the role of the enthalpy.
Acknowledgements.
[…]
Appendix A Equations of motion
In this section, we present the explicit expression for the equations of motion given in (3) together with the line element (6) for $\xi=\frac{1}{6}$ and $N(r)=1$, which are the following:
$\mathcal{E}_{\phantom{a}t}^{t}$ $=$ $2rN^{\theta}(N^{\theta})^{\prime}-\frac{1}{3}rN^{\theta}(N^{\theta})^{\prime}\phi^{2}+\frac{1}{2}r^{2}N^{\theta}(N^{\theta})^{\prime\prime}-\frac{1}{12}r^{2}N^{\theta}(N^{\theta})^{\prime\prime}\phi^{2}-\frac{1}{6}r^{2}\phi\phi^{\prime}N^{\theta}(N^{\theta})^{\prime}$
$+$ $\frac{f}{r^{2}}-\frac{f\phi^{2}}{6r^{2}}+\frac{f^{\prime}}{r}-\frac{f^{\prime}\phi^{2}}{r}+\frac{1}{4}r^{2}\big{(}(N^{\theta})^{\prime}\big{)}^{2}-\frac{1}{24}r^{2}\big{(}(N^{\theta})^{\prime}\big{)}^{2}\phi^{2}-\frac{1}{6}\phi\phi^{\prime}f^{\prime}+\frac{1}{6}(\phi^{\prime})^{2}f$
-- $\frac{1}{3}f\phi\phi^{\prime\prime}-\frac{2}{3r}\phi\phi^{\prime}f+U(\phi)+\Lambda=0,$
$\mathcal{E}^{t}_{\phantom{a}\theta}$ $=$ $2(N^{\theta})^{\prime}r-\frac{1}{3}r(N^{\theta})^{\prime}\phi^{2}+\frac{1}{2}r^{2}(N^{\theta})^{\prime\prime}-\frac{1}{12}r^{2}(N^{\theta})^{\prime\prime}\phi^{2}-\frac{1}{6}\phi\phi^{\prime}r^{2}((N^{\theta}))^{\prime}=0,$
$\mathcal{E}^{\theta}_{\phantom{a}t}$ $=$ $-\frac{1}{2}(N^{\theta})^{\prime\prime}f-\frac{1}{2}r^{2}(N^{\theta})^{\prime}(N^{\theta})^{2}+\frac{1}{12}(N^{\theta})^{\prime\prime}f\phi^{2}+\frac{1}{12}r^{2}(N^{\theta})^{\prime\prime}\phi^{2}(N^{\theta})^{2}+\frac{1}{2}N^{\theta}f^{\prime\prime}$
-- $\frac{1}{12}N^{\theta}f^{\prime\prime}\phi^{2}-r^{2}N^{\theta}\big{(}(N^{\theta})^{\prime}\big{)}^{2}+\frac{1}{6}r^{2}N^{\theta}\big{(}(N^{\theta})^{\prime}\big{)}^{2}\phi^{2}+\frac{1}{3r}(N^{\theta})^{\prime}f\phi^{2}+\frac{1}{3}r(N^{\theta})^{\prime}\phi^{2}(N^{\theta})^{2}$
-- $\frac{2(N^{\theta})^{\prime}f}{r}-2r(N^{\theta})^{2}(N^{\theta})^{\prime}+\frac{1}{6}(N^{\theta})^{\prime}\phi^{\prime}\phi f+\frac{1}{6}r^{2}(N^{\theta})^{\prime}\phi^{\prime}\phi(N^{\theta})^{2}+\frac{1}{3r}N^{\theta}\phi^{\prime}\phi f$
-- $\frac{1}{6}N^{\theta}\phi^{\prime}\phi f^{\prime}-\frac{N^{\theta}f}{r^{2}}+\frac{1}{6r^{2}}N^{\theta}f\phi^{2}=0,$
$\mathcal{E}^{r}_{\phantom{a}r}$ $=$ $-\frac{1}{2}(\phi^{\prime})^{2}f-\frac{2}{3r}\phi\phi^{\prime}f-\frac{1}{6}\phi\phi^{\prime}f^{\prime}-\frac{1}{6r}f^{\prime}\phi^{2}+\frac{f^{\prime}}{r}+\frac{1}{4}r^{2}\big{(}(N^{\theta})^{\prime}\big{)}^{2}-\frac{1}{24}r^{2}\big{(}(N^{\theta})^{\prime}\big{)}^{2}\phi^{2}$
-- $\frac{1}{6r^{2}}f\phi^{2}+\frac{f}{r^{2}}+U(\phi)+\Lambda=0,$
$\mathcal{E}^{\theta}_{\phantom{a}\theta}$ $=$ $\frac{1}{2}r^{2}N^{\theta}(N^{\theta})^{\prime\prime}+\frac{1}{12}r^{2}N^{\theta}(N^{\theta})^{\prime\prime}\phi^{2}+\frac{1}{2}f^{\prime\prime}-\frac{1}{12}f^{\prime\prime}\phi^{2}-\frac{1}{3}f\phi\phi^{\prime\prime}-\frac{3}{4}r^{2}\big{(}(N^{\theta})^{\prime}\big{)}^{2}$
$+$ $\frac{1}{8}r^{2}\big{(}(N^{\theta})^{\prime}\big{)}^{2}\phi^{2}+\frac{1}{3}rN^{\theta}(N^{\theta})^{\prime}\phi^{2}-2rN^{\theta}(N^{\theta})^{\prime}+\frac{1}{6}r^{2}N^{\theta}((N^{\theta})^{\prime})\phi^{\prime}\phi+\frac{1}{6}(\phi^{\prime})^{2}f$
-- $\frac{1}{3r}\phi\phi^{\prime}f-\frac{1}{3}\phi\phi^{\prime}f^{\prime}+\frac{f^{\prime}}{r}-\frac{1}{6r}f^{\prime}\phi^{2}+U(\phi)+\Lambda=0,$
$\mathcal{E}_{\phantom{a}\varphi}^{\varphi}$ $=$ $\frac{1}{2}f^{\prime\prime}-\frac{1}{12}f^{\prime\prime}\phi^{2}-\frac{1}{3}f\phi\phi^{\prime\prime}+\frac{1}{6}(\phi^{\prime})^{2}f-\frac{1}{3r}\phi\phi^{\prime}f-\frac{1}{3}\phi\phi^{\prime}f^{\prime}+\frac{f^{\prime}}{r}-\frac{1}{6r}f^{\prime}\phi^{2}$
-- $\frac{1}{4}r^{2}\big{(}(N^{\theta})^{\prime}\big{)}^{2}+\frac{1}{24}r^{2}\big{(}(N^{\theta})^{\prime}\big{)}^{2}\phi^{2}+U(\phi)+\Lambda=0,$
where, as before, (${}^{\prime})$ denotes the derivative with respect to coordinate $r$.